# Multiple Instance Verification

### v1

@@ B.3 Fact extraction and verification
For“fact extraction and verification” (FEVER), we collected the raw data of claims and evidence as in [33] and from the FEVER (2018) website https://fever.ai/dataset/fever.html (license see https://fever.ai/download/fever/license.html). The FEVER raw data was already split into training, validation, and test data, based on a pre-processed dump (June 2017) of Wikipedia pages—more precisely, the corresponding links are: https://fever.ai/download/fever/train.jsonl, https://fever.ai/download/fever/paper_dev.jsonl, https://fever.ai/download/fever/paper_test.jsonl, https://fever.ai/download/fever/wiki-pages.zip. For all three rounds of experiments, we used the full set of raw validation/test data (6616/6613 exemplars respectively) as our validation/test datasets. To construct our training dataset in each round of experiments, we randomly sampled 33,000 exemplars from the raw training set, with the proportion of the exemplar labels being approximately 50:50. In an exemplar, we truncated the amount of evidence in a bag, i.e., the bag size $N$, if it was greater than 47, and capped the number of tokens in a piece of evidence to 96.121212The truncations are purely for technical reasons (to limit the use of memory/computing), and we found it impacts less than 10% of the total samples.
For the feature extractor, we employed one from the Sentence-BERT (SBERT) family [28], called “multi-qa-MiniLM-L6-cos-v1”, initialized with weights pretrained based on 215M question-answer pairs from various sources. We also used SBERT’s native tokenizer to pre-process all our datasets. See more details, downloadable weights, and the tokenizer at https://www.sbert.net/docs/pretrained_models.html (Apache License 2.0, https://github.com/UKPLab/sentence-transformers/blob/master/LICENSE). We obtained the feature vector of any textual paragraph by putting it through the feature extractor, and retrieved the vector of SBERT’s special token “[CLS]”  [^13] Note the vector of SBERT’s special token “[CLS]” is different from that of the MSA benchmark’s “[CLS]”. from the models’ penultimate layer.
@@
Appendix C QMNIST: Results for larger average bag sizes
@@
(b) Quality of the explanations: average i-AUROC
@@
(d) Quality of the explanations: average i-AP
Figure 7: Test AUROC (6(a)), accuracy (6(c)), average i-AUROC (6(b)), average i-AP (6(d)) for QMNIST for varying numbers of training exemplars, from 8,000 to 24,000, with an average of 20 instances per target bag.
@@
(b) Quality of the explanations: average i-AUROC
@@
(d) Quality of the explanations: average i-AP
Figure 8: Test AUROC (7(a)), accuracy (7(c)), average i-AUROC (7(b)), average i-AP (7(d)) for QMNIST for varying numbers of training exemplars, from 8,000 to 24,000, with an average of 50 instances per target bag.
Learning curves for the QMNIST data for mean bag sizes 20 and 50 are shown in Figures 7 and 8 respectively, in the range of 8,000 to 24,000 training exemplars.
@@ Appendix D Does multi-head attention matter?
(b) Quality of the explanations
Figure 9: Model performance for the QMNIST task under our CAP framework (DBA and VEMA), with and without the multi-head linear projections. Left column (8(a)) is classification performance; right column (8(b)) shows quality of the explanations.
@@
(b) Quality of the explanations
Figure 10: Model performance for the signature verification task under our CAP framework (DBA and VEMA), with and without the multi-head linear projections. Left column (9(a)) is classification performance; right column (9(b)) shows quality of the explanations.
(a) Classification
(b) Quality of the explanations
Figure 11: Model performance for the FEVER task under our CAP framework (DBA and VEMA), with and without the multi-head linear projections. Left column (10(a)) is classification performance; right column (10(b)) shows quality of the explanations.
@@
Figures 9, 10, 11 report the out-of-sample results of the “multi-head” and “non-headed” versions of DBA and VEMA under the CAP framework, on the three tasks respectively.
Generally speaking, the differences between the multi-head and non-headed versions are small, either insignificant or marginally significant. For the CV tasks (QMNIST and signature verification), whilst multi-head VEMA obtains statistically indistinguishable improvements over the non-headed version, CAP-DBA has even slightly better performance by excluding the multi-head projection, albeit with little statistical significance. For the NLP task (FEVER), the results are inconclusive. Even though better quality of explanation is notable for models with multi-head projection, this nevertheless does not translate to better classification—actually, we obtain significantly worse classification performance—for both CAP-VEMA and CAP-DBA (cf. Figure 10(a)).
Appendix E Sensitivity of our models to different forms of LayerNorm
{NiceTabular}
[…]
@@ Appendix E Sensitivity of our models to different forms of LayerNorm
{NiceTabular}
[…]
@@
{NiceTabular}
[…]
@@
One alternative to pre-aggregation LayerNorm is the post-aggregation LayerNorm adopted in the transformer model, i.e. a standard LayerNorm without multi-headedness. We replace the pre-aggregation LayerNorm with this alternative in both CAP-VEMA and CAP-DBA, and compare performance before and after replacement.
The results, shown in Tables E, E, E for the three tasks respectively, are mixed, across classification performance and quality of the explanations, and across different tasks. The performance differences between the two treatments are generally insignificant or only marginally significant. The results indicate no clear winner and demonstrate that CAP-based models are not sensitive to different forms of LayerNorm when applied in the final step before output. Therefore, their performances are robust to the choices of output normalization.

### v2

@@ B.3 Fact extraction and verification
For“fact extraction and verification” (FEVER), we collected the raw data of claims and evidence as in Thorne et al. (2018) and from the FEVER (2018) website https://fever.ai/dataset/fever.html (license see https://fever.ai/download/fever/license.html). The FEVER raw data was already split into training, validation, and test data, based on a pre-processed dump (June 2017) of Wikipedia pages—more precisely, the corresponding links are: https://fever.ai/download/fever/train.jsonl, https://fever.ai/download/fever/paper_dev.jsonl, https://fever.ai/download/fever/paper_test.jsonl, https://fever.ai/download/fever/wiki-pages.zip. For all three rounds of experiments, we used the full set of raw validation/test data (6616/6613 exemplars respectively) as our validation/test datasets. To construct our training dataset in each round of experiments, we randomly sampled 33,000 exemplars from the raw training set, with the proportion of the exemplar labels being approximately 50:50. In an exemplar, we truncated the amount of evidence in a bag, i.e., the bag size $N$, if it was greater than 47, and capped the number of tokens in a piece of evidence to 96.161616The truncations are purely for technical reasons (to limit the use of memory/computing), and we found it impacts less than 10% of the total samples.
For the feature extractor, we employed one from the Sentence-BERT family (SBERT, Reimers and Gurevych (2019)), called “multi-qa-MiniLM-L6-cos-v1”, initialized with weights pretrained based on 215M question-answer pairs from various sources. We also used SBERT’s native tokenizer to pre-process all our datasets. See more details, downloadable weights, and the tokenizer at https://www.sbert.net/docs/pretrained_models.html (Apache License 2.0, https://github.com/UKPLab/sentence-transformers/blob/master/LICENSE). We obtained the feature vector of any textual paragraph by putting it through the feature extractor, and retrieved the vector of SBERT’s special token “[CLS]” [^17] Note the vector of SBERT’s special token “[CLS]” is different from that of the MSA benchmark’s “[CLS]”. from the models’ penultimate layer.
@@
(b) Key instance detection: average i-AUROC
@@
(d) Key instance detection: average i-AP
Figure 8: Test AUROC (8(a)), accuracy (8(c)), average i-AUROC (8(b)), average i-AP (8(d)) for QMNIST for varying numbers of training exemplars, from 8,000 to 24,000, with an average of 20 instances per target bag.
@@
(b) Key instance detection: average i-AUROC
@@
(d) Key instance detection: average i-AP
Figure 9: Test AUROC (9(a)), accuracy (9(c)), average i-AUROC (9(b)), average i-AP (9(d)) for QMNIST for varying numbers of training exemplars, from 8,000 to 24,000, with an average of 50 instances per target bag.
(a) Classification
(b) Key instance detection
Figure 10: Model performance for the QMNIST task under our CAP framework (DBA and VEMA), with and without the multi-head linear projections. Left column (10(a)) is classification performance; right column (10(b)) shows performance of detecting key instances.
Appendix C QMNIST: Results for larger average bag sizes
Learning curves for the QMNIST data for mean bag sizes 20 and 50 are shown in Figures 8 and 9 respectively, in the range of 8,000 to 24,000 training exemplars.
@@ Appendix D Does multi-head attention matter?
(b) Key instance detection
Figure 11: Model performance for the signature verification task under our CAP framework (DBA and VEMA), with and without the multi-head linear projections. Left column (11(a)) is classification performance; right column (11(b)) shows performance of detecting key instances.
@@
(b) Key instance detection
Figure 12: Model performance for the FEVER task under our CAP framework (DBA and VEMA), with and without the multi-head linear projections. Left column (12(a)) is classification performance; right column (12(b)) shows performance of detecting key instances.
@@
Figures 10, 11, 12 report the out-of-sample results of the “multi-head” and “non-headed” versions of DBA and VEMA under the CAP framework, on the three tasks respectively.
Generally speaking, the differences between the multi-head and non-headed versions are small, either insignificant or marginally significant. For the CV tasks (QMNIST and signature verification), whilst multi-head VEMA obtains statistically indistinguishable improvements over the non-headed version, CAP-DBA has even slightly better performance by excluding the multi-head projection, albeit with little statistical significance. For the NLP task (FEVER), the results are inconclusive. Even though better key instance detection performance is notable for models with multi-head projection, this nevertheless does not translate to better classification—actually, we obtain significantly worse classification performance—for both CAP-VEMA and CAP-DBA (cf. Figure 12(a)).
Layer Normalization Classification Key instance detection
@@
Layer Normalization Classification Key instance detection
@@
Layer Normalization Classification Key instance detection
@@
Appendix E Sensitivity of our models to different forms of LayerNorm
One alternative to pre-aggregation LayerNorm is the post-aggregation LayerNorm adopted in the transformer model, i.e., a standard LayerNorm without multi-headedness. We replace the pre-aggregation LayerNorm with this alternative in both CAP-VEMA and CAP-DBA, and compare performance before and after replacement.
The results, shown in Tables 13, 14, 15 for the three tasks, respectively, are mixed, across classification and key instance detection performance, and across different tasks. The performance differences between the two treatments are generally insignificant or marginally significant. The results indicate no clear winner and demonstrate that CAP-based models are not sensitive to different forms of LayerNorm when applied in the final step before output. Therefore, their performance is robust to the choice of output normalization.